OpenStudy (anonymous):

what is the slope of (0,-6) (4,6)

OpenStudy (anonymous):

\[m=\frac{ \Delta y }{ \Delta x }=\frac{ -6-6 }{ 4-0 }=-3\]

OpenStudy (anonymous):

how do you find the slope in 7th grade words please i have no idea on what to do

OpenStudy (anonymous):

my bad it should be 0-4 slope is +3

OpenStudy (anonymous):

Slope is \[\frac{ Y _{1}-Y _{2} }{X _{1}-X _{2} }\] so fill it in with the coordinates provided

OpenStudy (anonymous):

\[\frac{ -6 -6 }{0-4 }\] simplify -12/-4 = 3